# LLM-in-Sandbox Elicits General Agentic Intelligence

### v1

@@ 1 Introduction
In this work, we propose LLM-in-Sandbox—enabling LLMs to explore within a code sandbox—as a promising next step along this trajectory. As shown in Figure 1, the sandbox is essentially a virtual computer with terminal capabilities, widely used by code agents such as Claude Code (Anthropic, 2025a). While it is typically used for software engineering (Jimenez et al., 2023), we argue its potential extends far beyond coding. Computers are perhaps the most versatile platform ever created—virtually any task can be accomplished through them. This versatility stems from three meta-capabilities: external resource access (e.g., the internet), file management, and code execution. We hypothesize that combining LLMs with a virtual computer may unlock their potential for general intelligence.
@@ 2 LLM-in-Sandbox Elicits General Intelligence
To explore the full potential of this paradigm, our design of LLM-in-Sandbox emphasizes two principles: minimal—providing a basic code sandbox with these three capabilities, and exploratory—encouraging models to discover diverse solution strategies. In the following, we describe our sandbox environment (Section 2.1), the LLM-in-Sandbox workflow (Section 2.2), as well as experiments (Section 2.3) and analysis (Section 2.4) in general domains.
@@ 2.1 Code Sandbox
Within the code sandbox, we equip the model with three fundamental tools that together realize the core capabilities of a computer: (1) execute_bash for executing arbitrary terminal commands—the most fundamental yet versatile interface that enables virtually any computer operation, including but not limited to installing packages, managing files, and running programs; (2) str_replace_editor for file creation, viewing, and editing; and (3) submit for indicating task completion. Detailed specifications are provided in Appendix A.
@@ 2.3 Experiments on General Domains
Claude-Sonnet-4.5-Think 85.6 92.2 +6.6 56.9 63.3 +6.4 83.3 84.4 +1.1
GPT-5 87.8 97.9 +10.1 52.3 57.5 +5.2 81.1 81.6 +0.5
DeepSeek-V3.2-Thinking 89.8 97.7 +7.9 58.2 59.9 +1.7 76.7 77.8 +1.1
MiniMax-M2 71.3 76.3 +5.0 45.1 49.1 +4.0 54.0 68.4 +14.4
Kimi-K2-Thinking 90.2 94.4 +4.2 55.9 54.5 -1.4 74.4 77.6 +3.2
Qwen3-Coder-30B-A3B 17.9 42.1 +24.2 36.8 47.9 +11.1 50.2 55.8 +5.6
Qwen3-4B-Instruct-2507 41.3 35.4 -5.9 40.5 36.3 -4.2 56.2 50.7 -5.5
@@
Qwen3-Coder-30B-A3B 13.4 14.8 +1.4 27.3 24.0 -3.3 35.0 40.0 +5.0
Qwen3-4B-Instruct-2507 10.4 10.6 +0.2 30.8 5.8 -25.0 32.7 28.7 -4.0
@@ Results
As shown in Table 2, strong agentic models consistently benefit from LLM-in-Sandbox, with improvements observed across all evaluated domains—from computation-intensive tasks (Mathematics) to knowledge-intensive tasks (Chemistry, Biomedicine) to general capabilities (Instruction Following, Long-Context). The largest gains reach +24.2% (Qwen3-Coder on Mathematics). However, weaker models like Qwen3-4B-Instruct fail to benefit and even perform worse. We analyze the reasons in the following sections.
@@ 2.4 Analysis on Sandbox Utilization
To understand how models leverage the sandbox environment, we conduct a case study and quantitative analysis. Specifically, we focus on the three core capabilities of sandbox: external resource access, file management, and code execution. Code execution can serve many purposes; here we specifically track computation-oriented operations. We identify these behaviors through pattern matching on model actions: (1) external resources—network requests (e.g., curl, requests.get) and package installation (e.g., pip install); (2) file management—file I/O operations (e.g., open(), json.load) and shell commands (e.g., cat, grep); (3) computation—numerical solvers, iterative algorithms, and simulation loops. Detailed classification patterns are provided in Appendix D.
@@ 2.4.1 Case Study
In the chemistry task, the model is asked to predict molecular properties given only compound names. To solve this, the model autonomously installs Java runtime via apt-get and downloads the OPSIN library to convert chemical names to molecular structures—acquiring domain-specific tools unavailable in the base environment.
@@
In the instruction-following task, the model is asked to generate three sentences about medieval history, with a strict constraint: all sentences must have the same character count while using entirely different words. To satisfy this constraint, the model writes Python scripts to count characters, detect word overlaps, and iteratively refine candidate sentences—a task that would be extremely difficult through pure text generation.
@@ 2.4.2 Quantitative Analysis
Figure 2 shows strong models adapt their usage patterns to task requirements. Mathematics benefits most due to high computation frequency (43.4%)—models verify solutions through numerical computation. Chemistry shows the highest external resource frequency (18.4%) as models install domain-specific packages. In contrast, Biomedicine shows less stable improvements with the shortest exploration (6.5 turns), suggesting that models fail to fully leverage the sandbox environment for these tasks.
@@
Qwen-Coder 30.5 24.0
Qwen-4B 11.8 5.8
Average 35.6 48.9
@@
Table 4 compares sandbox utilization between strong and weak models. Strong models effectively leverage all three capabilities with high usage rate (6–21%), while the weak model (Qwen3-4B-Instruct) achieves far lower frequency (<<3%) despite taking nearly twice as many turns (23.7 vs. 12.6). This indicates that the weak model “wanders” in the sandbox without effective tool utilization—consuming more turns while accomplishing less.
@@ 3 LLM-in-Sandbox Reinforcement Learning Enhances Generalization
The preceding experiments demonstrate that the sandbox environment holds significant potential for enhancing general intelligence: strong agentic models consistently benefit across diverse domains. This raises a natural question: can we directly train LLMs within this paradigm to further unlock their potential? Motivated by this, we propose LLM-in-Sandbox Reinforcement Learning (LLM-in-Sandbox-RL), which trains models on general context-based tasks within the sandbox, enabling them to effectively explore the sandbox environment without requiring specialized agentic data.
@@ 3.1 Method
We propose training LLMs within sandbox environments configured with general-purpose data. Specifically, we adopt context-based tasks: each task consists of background materials (e.g., documents) and a task objective that must be completed based on these materials. Since completing the objective depends on the provided materials, models must actively explore the sandbox to find relevant information—naturally learning to leverage its capabilities. Meanwhile, our method uses only a general-purpose sandbox without task-specific configurations, making it easy to scale. Moreover, general-purpose tasks are much simpler to curate than software engineering tasks, enabling easy data scaling (Cheng et al., 2024; Jain et al., 2025). Overall, LLM-in-Sandbox-RL combines the benefits of sandbox-based training with general-domain data and easy scalability.
@@ 3.2 Experiments
Base LLM 41.3 35.4 40.5 36.3 56.2 50.7 10.4 10.6 30.8 5.8 32.7 28.7 11.2
LLM-RL 44.0 32.3 41.1 40.0 54.2 48.7 12.0 8.8 29.8 7.0 35.7 29.3 12.8
$\Delta$ +2.7 -3.1 +0.6 +3.7 -2.0 -2.0 +1.6 -1.8 -1.0 +1.2 +3.0 +0.6 +1.6
LLM-RL 47.9 50.2 46.5 47.7 56.9 59.8 10.0 14.4 35.0 16.8 33.7 37.7 12.4
$\Delta$ +6.6 +14.8 +6.0 +11.4 +0.7 +9.1 -0.4 +3.8 +4.2 +11.0 +1.0 +9.0 +1.2
@@
Base LLM 17.9 42.1 36.8 47.9 50.2 55.8 13.4 14.8 27.3 24.0 35.0 40.0 45.0
LLM-RL 14.6 42.1 39.9 46.0 47.6 56.7 17.0 17.2 30.8 21.8 34.3 26.3 47.6
$\Delta$ -3.3 0.0 +3.1 -1.9 -2.6 +0.9 +3.6 +2.4 +3.5 -2.2 -0.7 -13.7 +2.6
LLM-RL 17.3 43.5 40.2 49.1 51.3 56.9 16.4 18.4 27.8 30.5 34.0 42.7 48.0
$\Delta$ -0.6 +1.4 +3.4 +1.2 +1.1 +1.1 +3.0 +3.6 +0.5 +6.5 -1.0 +2.7 +3.0
@@
  • Inference Modes: Surprisingly, although trained exclusively in LLM-in-Sandbox mode, LLM-in-Sandbox-RL also improves LLM mode and even outperforms LLM-RL on most tasks, suggesting that agentic skills can transfer back to non-agentic generation. In contrast, LLM-RL primarily improves LLM mode, with limited gains in LLM-in-Sandbox mode.
@@
Base LLM 41.3 35.4 40.5 36.3 56.2 50.7 10.4 10.6 30.8 5.8 32.7 28.7 11.2
Math 43.1 49.0 43.1 46.8 55.3 61.8 10.6 12.8 28.5 14.3 34.3 36.7 15.2
SWE 46.9 30.0 46.2 42.5 53.8 51.1 10.4 12.2 29.3 7.8 33.3 31.7 17.4
Gen. in Prompt 45.4 33.1 46.9 46.6 56.0 60.2 11.8 13.2 28.0 12.8 32.3 29.0 16.2
Gen. in Sandbox 47.9 50.2 46.5 47.7 56.9 59.8 10.0 14.4 35.0 16.8 33.7 37.7 12.4
@@ 3.3 Analysis on Generalization
We analyze reasoning patterns in outputs of vanilla LLM mode before and after LLM-in-Sandbox-RL training. Specifically, we measure two categories of behaviors through pattern matching: (1) Structural Organization—Markdown formatting elements (headers, separators, bullet points, math blocks) that indicate explicit step-by-step reasoning; and (2) Verification Behaviors—phrases indicating self-checking (e.g., “let’s verify”, “check that”, confirmation markers). As shown in Table 9, both models exhibit increased structural organization and verification behaviors after training. These reasoning patterns, learned through multi-turn sandbox interaction where each action receives explicit feedback, transfer to LLM mode even without sandbox access.
@@ 5 LLM-in-Sandbox Goes Beyond Text Generation
  • File-Level Operations: Rather than describing what a file should contain, LLM-in-Sandbox directly produces actual files—.png, .mp4, .wav, .html—that users can immediately use, with grounded feedback from real execution.
  • Autonomous Tool Acquisition: Unlike predefined tool-use where LLMs call fixed APIs, LLM-in-Sandbox enables LLMs to autonomously discover, install, and learn to use arbitrary software libraries on demand—effectively granting unlimited tool access.
@@
Figure 5: LLM-in-Sandbox transcends the text-in-text-out paradigm. By granting LLMs access to a basic virtual computer, they can autonomously install tools, write and execute programs, and produce usable files—interactive webpages (.html), images (.png), videos (.mp4), and audio (.wav).
@@ Discussion.
Nevertheless, these cases reveal a promising direction: as LLMs become more capable and sandbox environments more sophisticated, LLM-in-Sandbox could evolve into a truly general-purpose digital creation system. We believe this paradigm—LLMs interacting with computational environments rather than generating text in isolation—represents a compelling path toward general intelligence.
@@ 6 Conclusion and Future Work
We introduce LLM-in-Sandbox, a paradigm that grants LLMs access to a virtual computer and show that strong LLMs exhibit emergent capabilities to leverage this environment for general tasks. Looking forward, we envision LLM-in-Sandbox becoming the default paradigm for serving LLMs: analytical tasks gain verifiable computation, long-context tasks benefit from file-based management, and creative tasks yield actual outputs—images, videos, applications—rather than text descriptions. We anticipate sandbox environments will become standard infrastructure, transforming LLMs from text generators into general-purpose digital workers.
@@
We propose LLM-in-Sandbox-RL, a lightweight RL method that trains sandbox interaction as a transferable skill using only general, non-agentic data. Looking ahead, we advocate for sandbox-native models where sandbox interaction becomes a first-class training objective—not only through large-scale RL with real environmental feedback, but also by incorporating sandbox-style reasoning into the pretraining stage itself.
@@ Appendix C Evaluation Details
UGPhysics is a comprehensive benchmark for evaluating physics problem-solving at the undergraduate level, spanning 13 core subjects. We sample 50 problems from each subject, yielding 650 problems in total. Responses are evaluated using an LLM-based judge (Zheng et al., 2023) with Qwen3-30B-A3B-Instruct-2507 (Yang et al., 2025a).

### v2

@@ 1 Introduction
In this work, we propose LLM-in-Sandbox—enabling LLMs to explore within a code sandbox—as a promising next step along this trajectory. As shown in Figure 1, the sandbox is essentially a virtual computer with terminal capabilities, widely used by code agents such as Claude Code (Anthropic, 2025a). While it is typically used for software engineering (Jimenez et al., 2023), we argue its potential extends far beyond coding. Computers are perhaps the most versatile platform ever created, virtually any task can be accomplished through them. This versatility stems from three meta-capabilities: external resource access (e.g., the internet), file management, and code execution. We hypothesize that combining LLMs with a virtual computer may unlock their potential for general intelligence.
@@ 2 LLM-in-Sandbox Elicits General Intelligence
To explore the full potential of this paradigm, our design of LLM-in-Sandbox emphasizes two principles: minimal, providing a basic code sandbox with these three capabilities, and exploratory, encouraging models to discover diverse solution strategies. In the following, we describe our sandbox environment (Section 2.1), the LLM-in-Sandbox workflow (Section 2.2), as well as experiments (Section 2.3) and analysis (Section 2.4) in general domains.
@@ 2.1 Code Sandbox
Within the code sandbox, we equip the model with three fundamental tools that together realize the core capabilities of a computer: (1) execute_bash for executing arbitrary terminal commands, (2) str_replace_editor for file creation, viewing, and editing, and (3) submit for indicating task completion. Specifically, execute_bash is a powerful meta-tool that enables models to install packages, run programs, and even programmatically create new tools on demand, thereby bootstrapping any additional functionality beyond the provided toolset. Detailed specifications are provided in Appendix A.
@@ 2.3 Experiments on General Domains
Claude-Sonnet-4.5-Think 85.6 92.2 +6.6 46.7 52.5 +5.8 83.3 84.4 +1.1
GPT-5 87.8 97.9 +10.1 41.0 45.8 +4.8 81.1 81.6 +0.5
DeepSeek-V3.2-Thinking 89.8 97.7 +7.9 37.6 45.2 +7.6 76.7 77.8 +1.1
MiniMax-M2 71.3 76.3 +5.0 30.3 35.8 +5.5 54.0 68.4 +14.4
Kimi-K2-Thinking 90.2 94.4 +4.2 40.6 42.3 +1.7 74.4 77.6 +3.2
Qwen3-Coder-30B-A3B 26.0 41.5 +15.5 23.9 28.9 +5.0 48.7 54.0 +5.3
Qwen3-4B-Instruct-2507 46.0 32.5 -13.5 22.9 25.6 +2.7 56.4 49.3 -7.1
@@
Qwen3-Coder-30B-A3B 14.4 18.2 +3.8 24.6 23.9 -0.7 34.7 40.3 +5.6
Qwen3-4B-Instruct-2507 10.2 9.4 -0.8 8.6 10.5 +1.9 33.7 29.0 -4.7
@@ Results
As shown in Table 2, strong agentic models consistently benefit from LLM-in-Sandbox, with improvements observed across all evaluated domains: from computation-intensive tasks (Mathematics) to knowledge-intensive tasks (Chemistry, Biomedicine) to general capabilities (Instruction Following, Long-Context). The largest gains reach +15.5% (Qwen3-Coder on Mathematics). However, weaker models like Qwen3-4B-Instruct fail to benefit and even perform worse. We analyze the reasons in the following sections.
@@ 2.4 Analysis on Sandbox Utilization
To understand how models leverage the sandbox environment, we conduct a case study and quantitative analysis. Specifically, we focus on the three core capabilities of sandbox: external resource access, file management, and code execution. Code execution can serve many purposes; here we specifically track computation-oriented operations. We identify these behaviors through pattern matching on model actions: (1) external resources: network requests (e.g., curl, requests.get) and package installation (e.g., pip install); (2) file management: file I/O operations (e.g., open(), json.load) and shell commands (e.g., cat, grep); (3) computation: numerical solvers, iterative algorithms, and simulation loops. Detailed classification patterns are provided in Appendix D.
@@ 2.4.1 Case Study
In the chemistry task, the model is asked to predict molecular properties given only compound names. To solve this, the model autonomously installs Java runtime via apt-get and downloads the OPSIN library to convert chemical names to molecular structures, acquiring domain-specific tools unavailable in the base environment.
@@
In the instruction-following task, the model is asked to generate three sentences about medieval history, with a strict constraint: all sentences must have the same character count while using entirely different words. To satisfy this constraint, the model writes Python scripts to count characters, detect word overlaps, and iteratively refine candidate sentences, a task that would be extremely difficult through pure text generation.
@@ 2.4.2 Quantitative Analysis
Figure 2 shows strong models adapt their usage patterns to task requirements. Mathematics benefits most due to high computation frequency (43.4%): models verify solutions through numerical computation. Chemistry shows the highest external resource frequency (18.4%) as models install domain-specific packages. In contrast, Biomedicine shows less stable improvements with the shortest exploration (6.5 turns), suggesting that models fail to fully leverage the sandbox environment for these tasks.
@@
Qwen-Coder 30.5 23.9
Qwen-4B 11.8 10.5
Average 35.6 49.6
@@
Table 4 compares sandbox utilization between strong and weak models. Strong models effectively leverage all three capabilities with high usage rate (6–21%), while the weak model (Qwen3-4B-Instruct) achieves far lower frequency (<<3%) despite taking nearly twice as many turns (23.7 vs. 12.6). This indicates that the weak model “wanders” in the sandbox without effective tool utilization, consuming more turns while accomplishing less.
@@ 3 LLM-in-Sandbox Reinforcement Learning Enhances Generalization
The preceding experiments demonstrate that the sandbox environment holds significant potential for enhancing general intelligence: strong agentic models consistently benefit across diverse domains. This raises a natural question: can we directly train LLMs within the sandbox to further unlock their potential? Motivated by this, we propose LLM-in-Sandbox Reinforcement Learning (LLM-in-Sandbox-RL), which trains models on general context-based tasks within the sandbox, enabling them to effectively explore the sandbox environment without requiring specialized agentic data.
@@ 3.1 Method
We propose training LLMs within sandbox environments configured with general-purpose data. Specifically, we adopt context-based tasks: each task consists of background materials (e.g., documents) and a task objective that must be completed based on these materials. Since completing the objective depends on the provided materials, models must actively explore the sandbox to find relevant information, naturally learning to leverage its capabilities. Meanwhile, our method uses only a general-purpose sandbox without task-specific configurations, making it easy to scale. Moreover, general-purpose tasks are much simpler to curate than software engineering tasks, enabling easy data scaling (Cheng et al., 2024; Jain et al., 2025). Overall, LLM-in-Sandbox-RL combines the benefits of sandbox-based training with general-domain data and easy scalability.
@@ 3.2 Experiments
Base LLM 46.0 32.5 22.9 25.6 56.4 49.3 10.2 9.4 8.6 10.5 33.7 29.0 11.2
LLM-RL 48.5 33.1 21.3 26.0 52.4 50.4 9.4 12.8 9.1 14.1 35.3 34.0 12.8
$\Delta$ +2.5 +0.6 -1.6 +0.4 -4.0 +1.1 -0.8 +3.4 +0.5 +3.6 +1.7 +5.0 +1.6
LLM-RL 51.0 53.3 23.0 30.3 57.8 63.3 10.8 13.2 10.6 17.6 33.7 38.7 12.8
$\Delta$ +5.0 +20.8 +0.1 +4.7 +1.3 +14.0 +0.6 +3.8 +2.0 +7.1 0.0 +9.7 +1.6
@@
Base LLM 26.0 41.5 23.9 28.9 48.7 54.0 14.4 18.2 24.6 23.9 34.7 40.3 45.0
LLM-RL 28.1 40.9 23.9 29.5 50.5 57.1 16.8 15.6 26.5 21.9 34.0 39.3 47.6
$\Delta$ +2.1 -0.6 0.0 +0.6 +1.8 +3.1 +2.4 -2.6 +1.9 -2.0 -0.7 -1.0 +2.6
LLM-RL 29.8 42.7 25.4 29.3 49.1 56.7 18.6 17.6 24.6 36.1 31.7 38.0 48.0
$\Delta$ +3.8 +1.2 +1.5 +0.4 +0.4 +2.7 +4.2 -0.6 0.0 +12.2 -3.0 -2.3 +3.0
@@
  • Inference Modes: Surprisingly, although trained exclusively in LLM-in-Sandbox mode, LLM-in-Sandbox-RL also improves LLM mode and even outperforms LLM-RL on most tasks, suggesting that agentic skills can transfer back to non-agentic generation.
@@
Base LLM 46.0 32.5 22.9 25.6 56.4 49.3 10.2 9.4 8.6 10.5 33.7 29.0 11.2
Math 53.5 51.0 23.4 31.4 58.2 60.0 9.6 13.6 9.5 11.3 33.0 37.0 15.2
SWE 47.5 36.5 21.3 21.2 57.8 47.8 9.4 9.2 11.6 9.9 33.0 33.0 17.4
Gen. in Prompt 47.7 37.7 24.0 26.4 58.0 58.4 9.6 16.6 10.2 15.1 33.0 34.3 16.2
Gen. in Sandbox 51.0 53.3 23.0 30.3 57.8 63.3 10.8 13.2 10.6 17.6 33.7 38.7 12.8
@@ 3.3 Analysis on Generalization
We analyze reasoning patterns in outputs of vanilla LLM mode before and after LLM-in-Sandbox-RL training. Specifically, we measure two categories of behaviors through pattern matching: (1) Structural Organization: Markdown formatting elements (headers, separators, bullet points, math blocks) that indicate explicit step-by-step reasoning; and (2) Verification Behaviors: phrases indicating self-checking (e.g., “let’s verify”, “check that”, confirmation markers). As shown in Table 9, both models exhibit increased structural organization and verification behaviors after training. These reasoning patterns, learned through multi-turn sandbox interaction where each action receives explicit feedback, transfer to LLM mode even without sandbox access.
@@ 5 LLM-in-Sandbox Goes Beyond Text Generation
  • File-Level Operations: Rather than describing what a file should contain, LLM-in-Sandbox directly produces actual files (e.g., .png, .mp4, .wav, .html)that users can immediately use, with grounded feedback from real execution.
  • Autonomous Tool Acquisition: Unlike predefined tool-use where LLMs call fixed APIs, LLM-in-Sandbox enables LLMs to autonomously discover, install, and learn to use arbitrary software libraries on demand, effectively granting unlimited tool access.
@@
Figure 5: LLM-in-Sandbox transcends the text-in-text-out paradigm. By granting LLMs access to a basic virtual computer, they can autonomously install tools, write and execute programs, and produce usable files: interactive webpages (.html), images (.png), videos (.mp4), and audio (.wav).
@@ Discussion.
Nevertheless, these cases reveal a promising direction: as LLMs become more capable and sandbox environments more sophisticated, LLM-in-Sandbox could evolve into a truly general-purpose digital creation system. We believe this paradigm, LLMs interacting with computational environments rather than generating text in isolation, represents a compelling path toward general intelligence.
@@ 6 Conclusion and Future Work
We introduce LLM-in-Sandbox, a paradigm that grants LLMs access to a virtual computer and show that strong LLMs exhibit emergent capabilities to leverage this environment for general tasks. Looking forward, we envision LLM-in-Sandbox becoming the default paradigm for serving LLMs: analytical tasks gain verifiable computation, long-context tasks benefit from file-based management, and creative tasks yield actual outputs (e.g., images, videos and applications) rather than text descriptions. We anticipate sandbox environments will become standard infrastructure, transforming LLMs from text generators into general-purpose digital workers.
@@
We propose LLM-in-Sandbox-RL, a lightweight RL method that trains sandbox interaction as a transferable skill using only general, non-agentic data. Looking ahead, we advocate for sandbox-native models where sandbox interaction becomes a first-class training objective, not only through large-scale RL with real environmental feedback, but also by incorporating sandbox-style reasoning into the pretraining stage itself.
@@ Appendix C Evaluation Details
UGPhysics is a comprehensive benchmark for evaluating physics problem-solving at the undergraduate level, spanning 13 core subjects. We sample 50 problems from each subject, yielding 650 problems in total. Responses are evaluated using an LLM-based judge (Zheng et al., 2023) with Qwen3-235B-A22B-Instruct-2507 (Yang et al., 2025a).